OpenStudy (anonymous):

When asked to write a point-slope equation for a line passing through the points (-3, -5) and (2, 4), a student wrote y-4=-9/5(x-2) . How can he fix this to write the correct equation?

OpenStudy (anonymous):

A). CHange the minus signtoplus signs B). change the sign of the slope C). this is correct do not change D). distibute -9/5 toeach term in the parentheses

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (sleepyhead314):

Hello @autumnmarietyler1 Sorry for the late reply :-/ Let us first see what slope the students was supposed to have slope = rise/run = (y2 - y1)/(x2 - x1) so the slope in this case would be 4 - (-5) 4 + 5 9 ---------- = --------- = ------ 2 - (-3) 2 + 3 5 which is a positive slope does the student have a positive slope as well? I hope this helped! If you have anymore questions feel free to tag me :) @sleepyhead314